Session: Feb 2002 (session 1)
Wavelength: 18. cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (X.Y. Huang at Sun Feb 10 01:42:28 MET 2002) Success

Medicina (G.Maccaferri at Sun Feb 10 01:57:26 MET 2002) Success

Urumqi ( at Sun Feb 10 03:57:40 MET 2002) Success

Torun (Kazimierz Borkowski at Mon Feb 11 09:51:58 MET 2002) Minor failures
No lock in BBC7 (so no useful data from it). No FS Tsys measurements.

Jodrell_Bank (Alastair Gunn at Tue Feb 12 11:15:00 MET 2002) Minor failures
Used Mk2 telescope. No data for first 5 minutes due to problems with LOs. Phase cal signal present only after 14:14 UT (day 39). Telescope taken offline at 04:30 UT (day 40) due to high winds. Back on line at 06:50 UT (day 40). FS crash at about 12:02 UT (day 40), restarted for new tape at 13:15, but data may be good until 12:40 UT. Several other minor problems noted in log.

Cambridge (Alastair Gunn at Tue Feb 12 11:15:42 MET 2002) Failure
Did not observe. Cambridge telescope link failed.

Westerbork (Tony Foley at Wed Feb 13 12:02:13 MET 2002) Minor failures
Lost scans on saturday 02:32 to 04:16 (mostly out of HA range anyway) and 19:12 to 19:45 (out of HA range). Other scans out of hour angle range replaced by local calibration. Some EMI -mostly low end of the band, BBC01 and BBC03

Effelsberg (A. Kraus at Wed Feb 13 19:03:49 MET 2002) Minor failures
High parity errors.

Onsala (John Conway at Wed Mar 20 14:41:12 CET 2002) Success